Two waves are described by y_1 = 6 cos 180t and y_

I know the answer is 3 rad/s, but I need help getting there. Please show your work.
Answer #1

here,

y1 = 6 * cos(180 * t)

y2 = 6 * cos(186 * t)

y1 + y2 = 6 * ( cos(180t) + cos(186t))

y1 + y2 = 6 * 2 * ( cos((186 + 180)t/2) * cos((186 - 180)t/2))

y1 + y2 = 6 * 2 * ( cos(366t/2) * cos((6t)/2))

y1 + y2 = 12* ( cos(183t) * cos(3 * t) )

so, 183 rad/s and 3 rad/s gives the same wave

so, the angular frequency of the resultant wave is 3 rad/s